You are building a cathode ray tube and you need to ensure that the electron gun emits electrons at 4500000m/s; if the field in the electron gun is created by two plates separated by 5.5cm, what voltage should you use?


1
Expert's answer
2021-03-16T09:01:34-0400

Explanations & Calculations


  • A simple way of the electron gun can be sketched as shown


  • What happens is that a force is generated on the electron ("\small F=Eq") which accelerates it towards the anode with the final velocity as described in the description.
  • Since the field is uniform, the acceleration is constant & to get a connection between the data, the four equations of motion can be employed.
  • Field between a plates is given by "\small E = \Large\frac{\Delta V}{d}"
  • Then,

"\qquad\qquad
\begin{aligned}
\small \rightarrow v^2&=\small u^2+2as\\
\small v^2 &=\small 0+2\cdot\frac{F}{m}\cdot d\\
&=\small 2\cdot\frac{\Delta Vq}{md}\cdot d\\
\small \Delta V&=\small \frac{mv^2}{2q}
\end{aligned}"

  • The same result can be gained by considering the work transferred on to the electron by the fiels

"\qquad\qquad
\begin{aligned}
\small \frac{1}{2}mv^2&=\small q\Delta V
\end{aligned}"

  • Mass of en electron is "\small 9.11\times10^{-31}kg" and its charge is "\small -1.6\times10^{-19}C"
  • Therefore,

"\qquad\qquad
\begin{aligned}
\small \Delta V &=\small \frac{9.11\times10^{-31}kg\times(4.5\times10^6ms^{-1})^2}{2\times1.6\times10^{-19}C}\\
&=\small \bold{57.649\,V}
\end{aligned}"